OpenStudy (fionacndg):

5 related algebra questions need help! will fan and medal please help,

OpenStudy (fionacndg):

Graph the coordinates of the specifics points in space your spacecraft will travel to. I have already done this... and they are Launch Area:___(1, 2)___ Point A:___(0, 3)___ Point B:___(-3, 0)___ Point C:___(-1, -4)___ You must show your work on each question below. Determine the equation of the line, in standard form, that will get your spacecraft from the Launch Area to Point A. Determine the equation of the line, in point-slope form, that will get your spacecraft from Point A to Point B. Determine the equation of the line, in slope-intercept form, that will get your spacecraft from Point B to Point C. Convert the equation you arrived at in question 2 into slope-intercept form. Make sure to include all of your work. Reflect back on this scenario and each equation you created. Would any restrictions apply to the domain and range of those equations? Explain your reasoning using complete sentences.

OpenStudy (fionacndg):

here is the graph i made

OpenStudy (fionacndg):

@greenglasses are you here?

OpenStudy (anonymous):

Sorry I was away.

OpenStudy (anonymous):

Just use the equation \[\frac{ y-y1 }{ x-x1}\] to find the slope of the lines, then plug that value (m) into the equation\[y- y1 = m(x-x1)\] to find the equation fo the line.

OpenStudy (fionacndg):

thanks! no problem!

OpenStudy (fionacndg):

Okay so for the first question, its 2-3/1-0 = -1/ 1 for m ?

OpenStudy (anonymous):

Yes. Alternately, you could always just read it off the graph, but knowing the equation is good in the long wrong.

OpenStudy (anonymous):

*run

OpenStudy (fionacndg):

okay great! i can do that for the rest of the problems, about the last... "Reflect back on this scenario and each equation you created. Would any restrictions apply to the domain and range of those equations? Explain your reasoning using complete sentences." What is it really talking about?

OpenStudy (anonymous):

Unlike a regular line, you want your lines to go from a certain point to another certain point, right?

OpenStudy (anonymous):

Wouldn't that mean you need to restrict the domain and range to make sure that you don't go past those certain points?

OpenStudy (fionacndg):

I think so... so I would need to restrict the domain and range so the "object" or point will not drift off to far? (pretending the point was a space object)

OpenStudy (anonymous):

Pretty much :)

OpenStudy (fionacndg):

Okay great! so for the first problem after I got m do I just plug the numbers from the coordinates back in and m and then i will get the answer from there?

OpenStudy (anonymous):

Yes, then rearrange to make it standard form. Oh, but you only have to plug in the numbers for one coordinate (the x1, y1 one).

OpenStudy (fionacndg):

ohh okay, so for 1 it would be y -2 = 1/-1(x-1) ? simplified?

OpenStudy (anonymous):

1 divided by -1 is just -1

OpenStudy (anonymous):

But yes.

OpenStudy (fionacndg):

great! i got y=−1x+3 thanks! i think I got it from here!!

OpenStudy (fionacndg):

woah y= -1 x +3

OpenStudy (anonymous):

Looks good to me. You can double check with this if you want https://www.desmos.com/calculator

OpenStudy (fionacndg):

thank you!

OpenStudy (anonymous):

Np.